# Homeomorphism theorem for sums of translates on the real axis

### v1

@@ 1. Introduction
where the field function $J$ is a function defined on $\mathbb{R}$, which is "admissible" for the kernels $K_{1},\ldots,K_{n}$ concave on $(-\infty,0)$ and on $(0,\infty)$ and having a singularity at $0.$
The sums of translates and the minimax problem for such functions were first considered by P. C. Fenton in 2000 [1]. He considered one kernel with assumptions of monotonicity, smoothness, singularity of its derivative at 0, and a concave field $J$ continuous at the ends of the segment. Fenton’s original goal was to prove P. D. Barry’s conjecture from 1962 on the growth of entire functions, which he succeeded in 1981 [2]. And although A. A. Goldberg proved this conjecture a little earlier [3], Fenton obtained other interesting results in the theory of entire functions using his approach [4], [5].
In our previous paper [6], we studied the minimax problem for the sums of translates on the real axis. The main method was a reduction to the minimax theorem proved by B. Farkas, B. Nagy and Sz. Gy. Révész for a segment (see [7], [8]). The uniqueness of the minimax point followed immediately from this reduction. To prove the uniqueness of the minimax point on the segment, the authors used the so-called homeomorphism theorem [9, Th. 7.1], which was also proved by them. Now, we prove a similar homeomorphism theorem for the real axis. As for the segment, this result provides the uniqueness of the minimax point on the real axis, independently of the specific reduction technique.
@@
Let $f$ be a function continuous on $[0,1]$ and $y_{0}=0<y_{1}<\ldots<y_{n}<1=y_{n+1}$ be interpolation nodes. Denote by $\pi_{n+1}$ the space of polynomials of degree at most $n+1$ and by $P_{\mathbf{y}}:C_{[0,1]}\to\pi_{n+1}$ the Lagrange interpolation operator
$P_{\mathbf{y}}f(t):=\sum\limits_{j=0}^{n+1}f(y_{j})\ell_{j}(t),\quad\ell_{j}(t):=\prod\limits_{i\neq j}\dfrac{t-y_{j}}{y_{i}-y_{j}}.$
It is easy to show that
@@
where $\|P_{\mathbf{y}}\|$ is the operator norm. Therefore, it is natural to minimize $\|P_{\mathbf{y}}\|$ by $\mathbf{y}$ to optimize the interpolation. It is known that $\|P_{\mathbf{y}}\|=\|\Lambda_{\mathbf{y}}\|_{C_{[0,1]}},$ where $\Lambda_{\mathbf{y}}(t):=\sum\limits_{j=0}^{n+1}|\ell_{j}(t)|$ is the Lebesgue function. Thus, this optimization problem is exactly a minimax problem for the functions $\Lambda_{\mathbf{y}}.$ Denote $\lambda_{j}(\mathbf{y}):=\max\limits_{t\in[y_{j},y_{j+1}]}\Lambda_{\mathbf{y}}(t),\ j=0,\ldots,n.$
@@
In 1977, T. A. Kilgore [11] proved Bernstein’s conjecture. Moreover, he showed the uniqueness of the equioscillation point. More precisely, Kilgore’s note [12] describing the proof of the statement “the minimax point is an equioscillation point” was first published, and a few months later, the complete proof was presented in [11]. Subsequently, C. R. de Boor and A. Pinkus [13] also proved Bernstein’s conjecture. In fact, they obtained the following general result.
@@
Let $S^{[0,1]}:=\{\mathbf{y}=(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}:\ 0<y_{1}<\ldots<y_{n}<1\}.$ Then the difference function
@@
$\min\limits_{0\leqslant j\leqslant n}\lim\limits_{y_{j+1}-y_{j}\to 0}\ \max\limits_{0\leqslant i\leqslant n-1}|\varphi_{i+1}(\mathbf{y})-\varphi_{i}(\mathbf{y})|=\infty$
@@
Shi dealt with the following minimax problem: find a vector $\mathbf{y}=(y_{1},\ldots,y_{n}),\ 0<y_{1}<\ldots<y_{n}<1,$ that minimizes $\max\limits_{j=0,\ldots,n}\varphi_{j}(\cdot).$ Under these assumptions, Shi proved that there exists a unique extremal point and it has the equioscillation property. Moreover, Shi obtained the homeomorphism theorem for the difference function
@@
In 2018, B. Farkas, B. Nagy and Sz. Gy. Révész presented a solution of the minimax problem and a homeomorphism result for sums of translates $F(\mathbf{y},t):=K_{0}(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j})$ on a torus [15]. Here $K_{0},\ldots,K_{n}:\ \mathbb{R}\to[-\infty,0)$ are $2\pi$-periodic functions, strictly concave on $(0,2\pi)$. Assuming that for each $j=0,\ldots,n$ the function $K_{j}$ belongs to $C^{2}(0,2\pi)$ with $K^{\prime\prime}_{j}<0$ and $K(0)=K(2\pi)=-\infty,$ they proved that the difference function of local maxima is a homeomorphism between its domain and $\mathbb{R}^{n}$. On the one hand, the sums of translates approach is more specific than Shi’s. On the other hand, the authors provided an example [15, Ex. 5.13] demonstrating that Shi’s result is not applicable in their settings.
In 2021, Farkas, Nagy and Révész proved a homeomorphism theorem for sums of translates on the segment [9, Th. 7.1]. This result is the basis of our proof, since to solve the problem on the real axis we reduce it to the case of the segment. For a precise formulation of the problems, we need a little more preparation.
Definition 1.1.
We extend $K$ by defining
If $K(0)=$-$\infty$, the kernel function $K$ is called singular.
When $K$ is defined on $(-\infty,0)\cup(0,\infty),$ the following condition, called generalized monotonicity, is important to us:
By the concavity of $K$, the set where $K^{\prime}$ is defined has full measure. We consider the limits of $K^{\prime}$ on this set.
Remark 1.1.
[…]
If $K$ is (strictly) decreasing on (p,0)(-p,0) and (strictly) increasing on $(0,p)$, then we call $K$ (strictly) monotone. Obviously, if $K$ is defined on $(-\infty,0)\cup(0,\infty)$ and monotone, then it satisfies ($GM$).
Remark 1.2.
There is a direct connection between the notions of monotonicity and generalized monotonicity ($GM$). Specifically, condition ($GM$) is equivalent to the existence of a number $c$ such that the kernel K(t)ctK(t)-ct is monotone.
Indeed, suppose that (GMGM) holds. Then, defining $c:=\lim\limits_{t\to\infty}K^{\prime}(t)$, we consider the function $\widetilde{K}(t):=K(t)-ct$. By (GMGM), its derivative satisfies
$\lim\limits_{t\to-\infty}\widetilde{K^{\prime}}(t)=\lim\limits_{t\to-\infty}K^{\prime}(t)-c\leqslant 0\quad\text{and}\quad\lim\limits_{t\to\infty}\widetilde{K^{\prime}}(t)=\lim\limits_{t\to\infty}K^{\prime}(t)-c=0.$
Since $K^{\prime}$ is non-decreasing at all points of its domain, this implies that $\widetilde{K^{\prime}}$ is monotone.
Hence, obviously, (GMGM) holds.
Definition 1.2.
Consider a segment $[a,b]$. In what follows, we denote by $\overline{S^{[a,b]}}$ the closed simplex
@@
The authors proved the homeomorphism theorem for a=0,b=1.a=0,\ b=1. We formulate this theorem for an arbitrary segment for convenience for our later application. The theorem on $[0,1]$ can be trivially extended to $[a,b]$ by applying the linear transformation
@@
that maps the segment $[a,b]$ onto $[0,1]$. Let us show this.
Let $K^{[a,b]}_{j},\ j=1,\ldots,n,$ be kernel functions defined on $(a-b,0)\cup(0,b-a)$ and $J^{[a,b]}$ be a field function on $[a,b]$. Assume that these kernels and field satisfy the conditions of Theorem B. Let us show that the homeomorphism theorem proven on $[0,1]$ implies Theorem B. Let
@@
We have for $t\in[a,b]$ that
(4) $K^{[a,b]}_{j}(t-y_{j})$ $=\mathscr{K}^{[0,1]}_{j}(\chi(t)-\chi(y_{j})),\quad j=1,\ldots,n,$
@@
Therefore, the difference function $\mathscr{D}^{[0,1]}$ is a homeomorphism between $\mathscr{R}^{[0,1]}$ and $\mathbb{R}^{n}.$ Denote $\boldsymbol{\chi}(\mathbf{y}):=(\chi(y_{1}),\ldots,\chi(y_{n}))$. Using (4), it is easy to see that
@@
In addition to [9, Th. 7.1], the authors obtained homeomorphism theorems with other conditions on the kernels and field. Conditions (3) may be replaced by so-called cusp conditions at the ends of the segment [9, Th. 7.5]. Moreover, if the derivatives of the kernel differences in (2) are bounded below by some $c>0,$ then the field can be arbitrary [9, Th. 2.1].
The minimax problem for the sums of translates on the segment was also deeply studied by Farkas, Nagy and Révész in [7], [8]. In particular, they proved that if the kernel $K^{[a,b]}$ is monotone, singular and strictly concave, then there exists a minimax point characterized by the equioscillation property. To prove the uniqueness of the equioscillation point, Farkas, Nagy and Révész apply Theorem B, which immediately implies this. Their research inspired the author to obtain similar results for the minimax problem on the real axis [6] and to write this paper.
@@
Let $K_{j}:(-\infty,0)\cup(0,\infty)\to\mathbb{R},\ j=1,\ldots,n,$ be kernels and $J:\mathbb{R}\to\underline{\mathbb{R}}$ be an $n$-field. We consider the “infinite closed simplex”
@@
(6) $D:R\to\mathbb{R}^{n},\ \mathbf{y}\mapsto(m_{1}(\mathbf{y})-m_{0}(\mathbf{y}),m_{2}(\mathbf{y})-m_{1}(\mathbf{y}),\ldots,m_{n}(\mathbf{y})-m_{n-1}(\mathbf{y})).$
@@
We prove the following result.
@@ 2. Simple lemmas about kernels and fields
In this section, we need the following extension of the fundamental theorem of calculus for concave functions. It is known [17, p. 9] that if $g$ is a concave function on an interval $I$, then for any $a,b]\leavevmode\nobreak\ \subset\leavevmode\nobreak\ I$
@@
  1. (1)
    Inequality (8) follows from definition of concavity. Its proof can be found e. g. in [18, Lemma 10].
@@
  2. (2)
    Sufficiently using (7), point 1 and (GMGM), we get
@@
    To get (10), one can group the terms in (9) with $t_{1}$ and $t_{2}$ together on the left respectively on the right hand side, then divide by $h$ and pass to the limit.
@@
1. Suppose that all $K^{\prime}_{j}$ be bounded below for large arguments. By our assumption and point 1 of Lemma 2.1, there are C,L>0C,\ L>0 such that
@@
Without loss of generality, assume that $\min\{t,t-y_{n}\}\geqslant L$. Using (7) and (11), we get
@@
2. Now assume that $K^{\prime}_{i}$ is not bounded below for some $i\in\{1,\ldots,n\}$ and for large $t.$ Denote
@@
Note that $f(\mathbf{y},\cdot)$ is concave on $(y_{n},\infty)$. Take arbitrary point $t_{0}\in(y_{n},\infty)$ where $f^{\prime}_{t}$ exists. For any $t\in(y_{n},\infty)$, we have [17, p. 12, Th. D]
@@
By point 1 of Lemma 2.1 for $K^{\prime}_{j},\ j=1\ldots,n,$ and our assumption regarding $K_{i}$, we conclude that $f^{\prime}_{t}(\mathbf{y},t_{0})<0$ for large $t_{0}$ as a sum of non-increasing functions and a function not bounded below. Due to (12), we obtain

### v2

@@ 1. Introduction
In this paper, we study sums of translates on the real axis. These functions generalize logarithms of weighted algebraic polynomials. Let $w\geqslant 0$ be a weight function and let $P(t):=(t-y_{1})\cdot\ldots\cdot(t-y_{n})$, $y_{1}\leqslant\ldots\leqslant y_{n}.$ Consider
$\log(w(t)|P(t)|)=\log w(t)+\sum\limits_{j=1}^{n}\log|t-y_{j}|.$
Replacing $\log|\cdot-y_{j}|$ by $K_{j}(\cdot-y_{j})$ and $\log w$ by $J$, we obtain the sum of translates function
$F(\mathbf{y},t):=J(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j}),\quad\mathbf{y}:=(y_{1},\ldots,y_{n}),\ y_{1}\leqslant\ldots\leqslant y_{n},\ t\in\mathbb{R},$
where $J:\mathbb{R}\to\underline{\mathbb{R}}:=\mathbb{R}\cup\{-\infty\}$ is the field function and the kernels $K_{1},\ldots,K_{n}$ are concave on $(-\infty,0)$ and on $(0,\infty)$ and have a singularity at $0.$
In our previous paper [6], we studied the minimax problem for the sums of translates on the real axis. The main method was a reduction to the minimax theorem for a segment proved by B. Farkas, B. Nagy and Sz. Gy. Révész (see [7], [8]). The uniqueness of the minimax point on the real axis followed immediately from this reduction. To prove the uniqueness of the minimax point on the segment, Farkas, Nagy and Révész used the so-called homeomorphism theorem [9, Th. 7.1], which was also proved by them. Now, we prove a similar homeomorphism theorem for the real axis. As for the segment, this result provides the uniqueness of the minimax point on the real axis, independently of the specific reduction technique.
Definition 1.1.
We say that the kernel $K$ is strictly concave if $K$ is strictly concave on (p,0)(-p,0) and on $(0,p)$.
If $K$ is (strictly) decreasing on (p,0)(-p,0) and (strictly) increasing on $(0,p)$, then $K$ is called (strictly) monotone.
We extend $K$ by defining
If $K(0)=$-$\infty$, the kernel function $K$ is called singular.
Definition 1.2.
@@
Let $f$ be a function continuous on $[0,1]$. Let us introduce the open simplex
$S^{[0,1]}:=\{\mathbf{y}:=(y_{1},\ldots,y_{n})\in\mathbb{R}^{n}:\ 0=:y_{0}<y_{1}<\ldots<y_{n}<y_{n+1}:=1\}.$
Denote by $\pi_{n+1}$ the space of polynomials of degree at most $n+1$ and by $P_{\mathbf{y}}:C_{[0,1]}\to\pi_{n+1}$ the Lagrange interpolation operator
$(P_{\mathbf{y}}f)(t):=\sum\limits_{j=0}^{n+1}f(y_{j})\ell_{j}(t),\quad\ell_{j}(t):=\prod\limits_{i=0,\ i\neq j}^{n+1}\dfrac{t-y_{i}}{y_{j}-y_{i}},\ j=0,\ldots,n+1.$
It is well-known, see [Rivlin, p. 88], that
@@
where $\|P_{\mathbf{y}}\|$ is the operator norm. Therefore, it is natural to minimize $\|P_{\mathbf{y}}\|$ by $\mathbf{y}$ to optimize the interpolation. It is known that $\|P_{\mathbf{y}}\|=\|\Lambda_{\mathbf{y}}\|_{C_{[0,1]}},$ where $\Lambda_{\mathbf{y}}(t):=\sum\limits_{j=0}^{n+1}|\ell_{j}(t)|$ is the Lebesgue function. Thus, this optimization problem is reduced to the minimax problem for the functions $\Lambda_{\mathbf{y}},$ i.e.,
$\min\limits_{\mathbf{y}}\|P_{\mathbf{y}}\|=\min\limits_{\mathbf{y}}\max\limits_{t\in[0,1]}|\Lambda_{\mathbf{y}}(t)|.$
Moreover, to obtain the maximum of $\Lambda_{\mathbf{y}}$ on $[0,1]$, it is sufficient to consider the local maxima $\lambda_{j}(\mathbf{y}):=\max\limits_{t\in[y_{j},y_{j+1}]}\Lambda_{\mathbf{y}}(t),\ j=0,\ldots,n.$
@@
In 1977, T. A. Kilgore [11] proved Bernstein’s conjecture. Moreover, he showed the uniqueness of the equioscillation point. More precisely, Kilgore’s note [12] describing the proof of the statement "the minimax point is an equioscillation point" was first published, and a few months later, the complete proof was presented in [11]. Subsequently, C. R. de Boor and A. Pinkus [13] also proved Bernstein’s conjecture by a different method. Their approach is valuable, as their observation can be applied in various other contexts. In fact, they obtained the following general result.
@@
The difference function
@@
$\lim\limits_{\min\limits_{0\leqslant j\leqslant n}(y_{j+1}-y_{j})\to 0}\ \max\limits_{0\leqslant i\leqslant n-1}|\varphi_{i+1}(\mathbf{y})-\varphi_{i}(\mathbf{y})|=\infty$
@@
Shi dealt with the following minimax problem: find a vector $\mathbf{y}=(y_{1},\ldots,y_{n})$ with $0<y_{1}<\ldots<y_{n}<1,$ that minimizes $\max\limits_{j=0,\ldots,n}\varphi_{j}(\cdot).$ Under these assumptions, Shi proved that there exists a unique extremal point and it has the equioscillation property. Moreover, Shi obtained the homeomorphism theorem for the difference function
@@
The sums of translates and the minimax problem for such functions were first considered by P. C. Fenton in 2000 [1]. He considered one kernel with assumptions of monotonicity, smoothness, singularity of its derivative at 0, and a concave field $J$ continuous at the ends of the segment. Fenton’s original goal was to prove P. D. Barry’s conjecture from 1962 on the growth of entire functions, which he succeeded in 1981 [2]. More precisely, A. A. Goldberg proved this conjecture a little earlier [3], but Fenton obtained other interesting results in the theory of entire functions using his approach [4], [5].
In 2018, B. Farkas, B. Nagy and Sz. Gy. Révész presented a solution of the minimax problem and a homeomorphism result for sums of translates $F(\mathbf{y},t):=K_{0}(t)+\sum\limits_{j=1}^{n}K_{j}(t-y_{j})$ on a torus [15]. Here $K_{0},\ldots,K_{n}:\ \mathbb{R}\to[-\infty,0)$ are $2\pi$-periodic functions, strictly concave on $(0,2\pi)$. Assuming that for each $j=0,\ldots,n$ the function $K_{j}$ belongs to $C^{2}(0,2\pi)$ with $K^{\prime\prime}_{j}<0$ and $K_{j}(0)=K_{j}(2\pi)=-\infty,$ they proved that the difference function of local maxima is a homeomorphism between its domain and $\mathbb{R}^{n}$. On the one hand, the sums of translates approach is more specific than Shi’s. On the other hand, the authors provided an example [15, Ex. 5.13] demonstrating that Shi’s result is not applicable in their settings.
In 2021, Farkas, Nagy and Révész proved a homeomorphism theorem for sums of translates on the segment [9, Th. 7.1]. The minimax problem for the sums of translates on the segment was also deeply studied by Farkas, Nagy and Révész in [7], [8]. In particular, they consider positive numbers $r_{1},\ldots,r_{n}$, a kernel $K^{[a,b]}$ and sums of translates of the following form
$F(\mathbf{y},t)=J(t)+\sum\limits_{j=1}^{n}r_{j}K^{[a,b]}(t-y_{j}),\quad\mathbf{y}\in\overline{S^{[a,b]}},\ t\in[a,b],$
where
@@
They proved that if the kernel $K^{[a,b]}$ is monotone, singular and strictly concave, then there exists a minimax point characterized by the equioscillation property. To prove the uniqueness of the equioscillation point, Farkas, Nagy and Révész apply the homeomorphism theorem. Their research inspired the author to obtain similar results for the minimax problem on the real axis [6] and to write this paper.
@@
The authors proved the homeomorphism theorem for a=0,b=1.a=0,\ b=1. For convenience of further application, we formulate this theorem for an arbitrary segment. The theorem on $[0,1]$ can be trivially extended to $[a,b]$ by applying the linear transformation
@@
that maps the segment $[a,b]$ onto $[0,1]$. Let us show that the homeomorphism theorem proven on $[0,1]$ implies Theorem B.
Assume that kernels $K^{[a,b]}_{j},\ j=1,\ldots,n,$ and a field $J^{[a,b]}$ satisfy the conditions of Theorem B. Let
@@
We have for $t,c\in[a,b]$ that
(4) $K^{[a,b]}_{j}(t-c)$ $=\mathscr{K}^{[0,1]}_{j}(\chi(t)-\chi(c)),\quad j=1,\ldots,n,$
@@
Therefore, the difference function $\mathscr{D}^{[0,1]}$ is a homeomorphism between the regularity set $\mathscr{R}^{[0,1]}$ and $\mathbb{R}^{n}.$ Denote $\boldsymbol{\chi}(\mathbf{y}):=(\chi(y_{1}),\ldots,\chi(y_{n}))$. Using (4), it is easy to see that
@@
In addition to [9, Th. 7.1], the authors obtained homeomorphism theorems with other conditions on the kernels and field. Conditions (3) of Theorem B may be replaced by so-called cusp conditions at the ends of the segment [9, Th. 7.5]. Moreover, if the derivatives of the kernel differences in (2) are bounded below by some $c>0,$ then the field can be arbitrary [9, Th. 2.1].
@@
By the concavity of a kernel $K$, the set where $K^{\prime}$ is defined has full measure. In the following, we consider $K^{\prime}$ and the limits of $K^{\prime}$ on this set. When $K$ is defined on $(-\infty,0)\cup(0,\infty),$ the following condition, called generalized monotonicity, is important to us:
Remark 1.2.
Note that if (GMGM) holds, then $\lim\limits_{t\to-\infty}K^{\prime}(t)$ and $\lim\limits_{t\to\infty}K^{\prime}(t)$ are finite. Indeed, since $K^{\prime}$ is non-increasing at all points of its domain, the limits in (GMGM) exist in the extended sense (taking values in the extended real line), and $\infty<\lim\limits_{t\to-\infty}K^{\prime}(t),\ \lim\limits_{t\to\infty}K^{\prime}(t)<\infty.$ Now the finiteness of these limits immediately follows from the property (GMGM).
Remark 1.3.
Obviously, if $K$ is defined on $(-\infty,0)\cup(0,\infty)$ and monotone, then it satisfies (GMGM).
On the other hand, condition ($GM$) for a kernel function $K$ is equivalent to the existence of a number $c$ such that the kernel K(t)ctK(t)-ct is monotone. Indeed, suppose that ($GM$) holds. Consider the kernel $\widetilde{K}(t):=K(t)-ct$, where $c:=\lim\limits_{t\to\infty}K^{\prime}(t)$. We get
$\lim\limits_{t\to\infty}\widetilde{K^{\prime}}(t)=\lim\limits_{t\to\infty}K^{\prime}(t)-c=0.$
Comparing this with property ($GM$), one can also conclude that
$\lim\limits_{t\to-\infty}\widetilde{K^{\prime}}(t)=\lim\limits_{t\to-\infty}K^{\prime}(t)-c\leqslant\lim\limits_{t\to\infty}K^{\prime}(t)-c=0.$
Due to the concavity of $K$, the function $\widetilde{K^{\prime}}$ is non-increasing at all points of its domain, therefore $\widetilde{K^{\prime}}(t)\leqslant 0$ almost everywhere on $(-\infty,0)$, and hence $\widetilde{K}$ is non-increasing on $(-\infty,0)$. Similarly, $\lim\limits_{t\to\infty}\widetilde{K^{\prime}}(t)=0$ implies that $\widetilde{K^{\prime}}\geqslant 0$ a.e. on $(0,\infty)$, that guarantees that $\widetilde{K}$ is non-decreasing on $(0,\infty)$. Consequently, the kernel $\widetilde{K}$ is monotone.
Hence, obviously, (GMGM) holds.
Let us introduce the "infinite closed simplex"
@@
(6) $D:R\to\mathbb{R}^{n},\quad\mathbf{y}\mapsto(m_{1}(\mathbf{y})-m_{0}(\mathbf{y}),m_{2}(\mathbf{y})-m_{1}(\mathbf{y}),\ldots,m_{n}(\mathbf{y})-m_{n-1}(\mathbf{y})).$
@@
The main result of this paper is the following.
@@
The proof of Theorem 1.1 is organized as follows. In Sections 2 and 3, we prove auxiliary results concerning kernels and fields. Section 4 is devoted to showing that $D$ is a local homeomorphism, while in Section 5 we prove that $D$ is a proper map. Finally, in Section 6, we apply Ho’s theorem, which states that a proper local homeomorphism is global.
@@ 2. Simple lemmas about kernels and fields
In this section, we need the following extension of the fundamental theorem of calculus for concave functions. It is known [17, p. 9] that if $g$ is a concave function on an interval $I$, then for any $a,b]\penalty 10000\ \subset\penalty 10000\ I$
@@
  1. label=()
    Inequality (8) follows from the definition of concavity. Its proof can be found, e. g., in [18, Lemma 10].
@@
  2. lbbel=()
    Sufficiently using (7), Part (a) and the property (GMGM) for $K$, we get
@@
    To get (10), one can separate the terms in (9) with respect to $t_{1}$ and $t_{2}$, placing them on the left- and right-hand sides respectively, then divide by $h$ and take the limit $h\to 0$.
@@
1. Let $y_{i}:=\max\limits_{1\leqslant j\leqslant n}\{y_{j}\}$. Suppose that all $K^{\prime}_{j}$ be bounded below for large arguments. By our assumption and Part (a) of Lemma 2.1, there are C,L>0C,\ L>0 such that
@@
We consider $t\geqslant L$ such that $\min\{t,t-y_{i}\}\geqslant L$. Using (7) and (11), we get
@@
2. Now assume that $K^{\prime}_{s}$ is not bounded below for some $s\in\{1,\ldots,n\}$ and for large $t.$ Denote
@@
Note that $f(\mathbf{y},\cdot)$ is concave on $(y_{i},\infty)$. Take arbitrary point $t_{0}\in(y_{i},\infty)$ where $f^{\prime}_{t}$ exists. For any $t\in(y_{i},\infty)$, we have [17, p. 12, Th. D]
@@
By Part (a) of Lemma 2.1 for $K^{\prime}_{j},\ j=1\ldots,n,$ and our assumption regarding $K_{s}$, we conclude that $f^{\prime}_{t}(\mathbf{y},t_{0})<0$ for large $t_{0}$ as a sum of non-increasing functions and a function not bounded below. Due to (12), we obtain